### v1

@@ Abstract
Using logistic regression, we developed a classification model that predicts driver yielding based on the identified variables. Our analysis indicates that vehicle speed, the presence of parking lots, proximity to parks or schools, and the width of major road crossings significantly influence driver yielding at unsignalized intersections. This study contributes to one of the most comprehensive driver-pedestrian datasets in the US, offering valuable insights for traffic safety improvements. By making this information available, our study will support communities across Minnesota and the United States in their ongoing efforts to improve road safety for pedestrians.
@@ 1 Introduction
2 Literature Review
The literature is divided into four main sections. The first section considers driver-yielding behavior at the intersection, the second section discusses research on pedestrian behavior, the third section is on intersection characteristics, and the last section examines relevant and related crash literature.
2.1 Driver Behavior and Characteristics
The investigation into factors influencing driver yielding at intersections, including both signalized and unsignalized ones as well as mid-block crossings, has been a significant area of focus in recent studies [11, 12, 13, 14, 8]. These studies have identified multiple determinants of driver yielding, with emphasis on the crossing environment. Key variables include vehicle speed, driver assertiveness, and vehicular volume. Research efforts have also been directed towards developing various models to better understand driver yielding behavior, such as critical gap models and logistic regression models [6, 15, 7]. These models have highlighted essential factors like gap size, the influence of vehicles in the opposite direction yielding, and the type of vehicle involved. Notably, studies by Schroeder et al. [15] have established a strong link between assertive pedestrian behavior and an increased likelihood of driver yielding, as well as the significant impact that vehicle platoons have in reducing driver yielding. A comprehensive synthesis of existing literature reveals several common factors influencing driver-yielding behavior. Vehicle speed is consistently identified as a crucial factor [16, 15, 17, 18]. The presence of a following vehicle [15, 19], the yielding decisions of drivers in the opposite direction [6, 15], the type of vehicle [19], and the vehicle’s driving direction [20] all significantly contribute to shaping driver behavior at intersections.
2.2 Pedestrian Behavior and Characteristics
@@
In addition to driver behavior, pedestrian behavior at intersections can influence the drivers’ yielding behavior. For example, there is a methodology for analyzing unsignalized intersections based on gap acceptance behavior [21], and gap acceptance theory is widely used to evaluate the safe operation of the road at uncontrolled intersections [22]. Pedestrians have to wait for safe gaps to cross at intersections. This is a high risk for pedestrians because the wrong judgment of the accepted gap can result in traffic crashes. From the perspective of a pedestrian, a pedestrian will either accept or reject the gap in the situation of vehicle-pedestrian interaction, and this will force the driver to react differently if the driver accepts a shorter gap. A summary of work studying such pedestrian behavior is provided below.
Studies have investigated various aspects of pedestrian behavior, from temporal and spatial gap acceptance to influencing factors like gender, group size, and carrying baggage [22, 23, 24, 8]. For instance, spatial gaps selected by pedestrians follow a lognormal distribution and acceptance probability decreases with conflicting vehicle speed [22]. Encouraging walking or biking has also been found to reduce the total number of crashes, indicating that increased pedestrian mobility does not necessarily lead to more crashes [25]. Compared with pedestrians at a marked crosswalk, pedestrians are less likely to accept driver’s yielding, and drivers decelerate more for pedestrians in a crosswalk than pedestrians outside the crosswalk. For pedestrians, the waiting time positively correlates with the number of lanes and vehicle volume. Collectively, these insights contribute to a nuanced understanding of pedestrian dynamics at intersections, highlighting areas for potential intervention and urban planning.
2.3 Roadway and Traffic Characteristics
Intersection characteristics are the factors that go beyond individual driver or pedestrian behavior. The layout and design of the intersection have been the subject of several studies. The idea is that certain design features may be more or less welcoming to pedestrians and thus make transportation infrastructure safer (or less safe) for pedestrians. Consequently, intersection design has been the subject of several studies that attempt to identify intersection features that correlate positively to improvements in pedestrian safety and thus can be used to determine guidelines for the construction of future intersections or modifications of existing ones.
Research on the intersection angle has demonstrated that left-turning drivers’ yielding behavior varies depending on the angle, with drivers at obtuse-angled intersections showing more non-yielding behavior, and those at acute-angled intersections yielding more frequently [26, 20]. The geometric characteristics of the intersection also influence vehicle speed and collision possibilities. Engineering treatments such as signalization and crossing marks have been shown to have a significant impact on yielding behavior. Red signal or beacon treatments and pedestrian-activated signals have been found effective in improving compliance rates [27, 28]. Road signs that prompt drivers to yield have also been effective in increasing safe yielding distances, although activation rates of pedestrian lights vary with traffic conditions [29]. Overall, intersection characteristics present a multifaceted area of study with implications for safety enhancements through design adjustments and engineering treatments.
2.4 Crashes at Intersections
Research on pedestrian crashes at intersections, which often result from driver non-yielding, has evolved significantly over time. Early studies, like Herms’ analysis of 400 intersections, revealed higher pedestrian crash rates in marked crosswalks, albeit without accounting for pedestrian volume differences [30]. Zegeer et al. later found that on two-lane roads, marked crosswalks did not significantly change crash rates compared to unmarked ones [31], indicating the role of other factors in pedestrian safety.
Advancements in crash prediction, such as the neural network model by Akin and Akba [19], have improved our understanding of the factors influencing intersection crashes. This model, with 90.9% accuracy in Macomb County, Michigan, underscores the importance of specific interaction variables in crash outcomes.
Sociological studies, like Morency et al.’s work [32], link pedestrian crashes to broader social issues, such as economic disparities. Detailed analyses, such as in Alameda County, California, highlight the complex interplay between pedestrian and vehicle volumes, intersection design, and crash risks [33].
Overall, while there are commonalities between vehicle crash research and driver-yielding studies, crash analyses often incorporate a wider array of factors, though significant overlaps exist.
2.5 Summary of the Literature
Based on the literature reviewed, there are several features, both of drivers and of the infrastructure, that lead to higher-yielding rates. A summary of important features and the relevant literature is provided in Table 1.
Table 1: Variables Influence the Yielding Behavior.
Event Features Site Features
Feature Name Literature Feature Name Literature
Time [19], [34] Posted Speed [16]
Party Size [13], [6], [15], [8] Number of Lanes [13], [17]
Pedestrian Position [13], [18] Signage [6], [15], [27]
Pedestrian Type [23], [17], [35],[8] Markings [19]
Vehicle Speed [16], [15], [17], [18] Land Use [33]
Opposite Direction [6], [15] Lighting [29], [19]
Following Vehicles [15], [19] Road Surface [16], [19]
Vehicle Type [15], [19],[24] Bus Stops [2]
Type of Yield [6] Traffic Volume [16], [33]
Weather [19] Bike Facility [36], [37], [38]
Vehicle Turning Direction [20] Intersection Angle [26]
Pedestrian gender [8], [24] Lane width [39]
Vehicle speed stands out in most studies as the primary determinant of driver-pedestrian interaction outcomes. Techniques such as radar and computer vision technologies have been employed to measure vehicle speed. Additionally, the driver’s yielding decision is also influenced by the presence of other vehicles, such as the following vehicle, and whether the opposite direction vehicle yield and the driving direction and the type of vehicle also have effects. From the pedestrian’s perspective, the pedestrian volume, party size, gender, and type of pedestrian are factors that are identified in the literature as important. The time and weather of the event are also documented in some studies.
Site-specific features also play a vital role in driver-yielding decisions. Signs, markings, built environment, land use, nearby bus stops, and demographic characteristics are often cited as essential variables. Traffic volume is particularly highlighted as one of the most critical factors for traffic safety [39].
At the macroscopic level, numerous studies emphasize the positive influence of increased walking or biking on overall traffic safety and public health [40, 41, 37, 42]. The development of better infrastructure for pedestrians and cyclists is seen as a pathway to a sustainable future, reinforcing the idea that efforts to enhance pedestrian safety are likely to result in additional benefits. Safer roads promote increased pedestrian activity, thereby creating a cycle that further contributes to improved safety for all road users.
@@ 3 Data Collection and Description
One of the contributions of this work is the collection of a large-scale, open-source naturalistic dataset for driver yielding. We collected over 3,300 driver-pedestrian interactions at 18 different unsignalized intersections. The dataset, which focuses on driver-pedestrian interactions, is unique in both its scale and the depth of information it provides. In this section, we first describe a summary of the selected sites and the data collection technique, followed by a summary of the selected sites, and finally present summary statistics of the collected data. It is important to note that the sites are anonymized in our reporting to maintain privacy standards.
@@ 3.1 Site Characteristics
The selection of sites was guided by a set of criteria derived from the literature reviewed in Section 2. These criteria include factors like traffic volume, pedestrian density, and urban setting. Our chosen sites, spread across diverse geographical locations, provide a representative cross-section of various urban traffic scenarios.
The table below summarizes the characteristics of each site, including the number of lanes, posted speed limits, type of road markings, AADT, and intersection geometry. This diversity in site characteristics allows for a robust analysis of driver-pedestrian interactions under different traffic conditions.
Table 2: Site Characteristics.
[b]
Site Num. of Lanes (Main Street) Posted Speed Markings AADT Shape
Site 1 4 35 Unmarked 14600 T-shape
Site 2 2 35 Unmarked 14800 T-shape
Site 3 2 30 Standard1 14800 T-shape
Site 4 2 30 Unmarked 10700 four-way
Site 5 2 30 Unmarked 10500 four-way
Site 6 3 35 Continental 6200 four-way
Site 7 4 35 Continental 3300 T-shape
Site 8 4 30 Unmarked 10200 four-way
Site 9 3 35 Unmarked 5400 four-way
Site 10 4 35 Unmarked 18800 four-way
Site 11 3 30 Unmarked 18100 four-way
Site 12 2 30 Unmarked 10300 four-way
Site 13 2 35 Continental 7400 four-way
Site 14 2 30 Standard 7100 four-way
Site 15 3 30 Unmarked 16400 four-way
Site 16 2 30 Standard 12100 four-way
Site 17 4 30 Unmarked 12800 T-shape
Site 18 4 30 Unmarked 12500 four-way
@@ 3.2 Data Collection and Cleaning
To investigate driver-yielding behavior, we utilized the traffic information monitor (TIM) platform developed by the Minnesota Traffic Observatory (MTO) at the University of Minnesota. The TIM is a customized sensor pack that consists of a pole-mounted video camera that can extend up to 30 feet, 12 V batteries that can sustain up to four weeks of operation, a Raspberry Pi computer, and a watertight metal enclosure to protect the batteries and computer from the elements. In this study, we used the TIM and the camera to capture video data of pedestrians and drivers approaching intersections in a naturalistic setting, as depicted in Fig. 1.
@@
In the summer of 2021, we strategically deployed TIMs across 18 diverse sites in Minnesota. These sites were chosen to encapsulate a broad spectrum of pedestrian and driver interactions, environmental contexts, and traffic patterns. The TIMs recorded extensive video data during daylight hours over a period of one to two weeks at each site, focusing on capturing naturalistic driver and pedestrian behavior at intersections.
@@
We collect more than 50 features for each interaction, and each pair of interactions needs to be collected individually. To extract the driver-pedestrian interactions accurately and analyze the data efficiently, a trained data collector manually extracted individual interaction events of all the variables. The speed extraction methodology computes speed by measuring the time each vehicle takes to pass two known positions at the intersection. Building upon our preliminary work cited in [10], which involved data from three sites, this study expands the dataset to include 18 locations. It is noteworthy that the data collection coincided with the pandemic period, adding a unique dimension to the study.
@@
In this study, a “codable driver-pedestrian interaction” is characterized as an event where either the driver yields to the pedestrian or the pedestrian forces to yield for giving the right of way to the vehicle. A ”close call” is defined as a near-miss incident in which an imminent collision between a vehicle and a pedestrian is narrowly averted, often necessitating quick evasive actions by either or both parties involved. It’s worth noting that this measure is somewhat subjective, especially given that data collection was performed by a single individual. The term ”number of bus stops” refers to the count of bus stops located within one block of the main road intersecting the observed location. For further details on the variables collected and their implications, we direct readers to our publicly available repository and to Tables 3.1 and 3 in the following sections.
3.3 Summary of Collected Data
In total, the data consists of 3,314 interactions from the 18 unsignalized intersections. To our knowledge, this represents the largest dataset of its kind, with each interaction containing more than 50 variables. The collected data can be found at our published repository [43]  [^1] https://github.com/tianyi17/pedestrian_yielding_data_MN. No collisions were observed in the dataset. Variations in driver-yielding and non-yielding events were observed among the sites, as illustrated in Fig. 2. Sites 3 and 16 displayed a higher total number of interactions than any other site, along with a larger proportion of drivers yielding compared to drivers non-yielding. In contrast, sites 4 and 5, located a block apart on the same corridor and situated along a park, presented a unique scenario. Only site 5 features a bus stop, leading to a higher total number of interactions (ranking third in volume among the 18 sites), which seems intuitive. However, the ratio of drivers yielding to drivers non-yielding was notably low at both sites. Interestingly, site 5 deviates from the general trend of increased interactions correlating with improved driver-yielding behavior.

### v2

@@ Abstract
Using logistic regression, we developed a classification model that predicts driver yielding based on the identified variables. Our analysis indicates that vehicle speed, the presence of parking lots, proximity to parks or schools, and the width of major road crossings significantly influence driver yielding at unsignalized intersections. Through our findings and by publishing one of the most comprehensive driver-pedestrian datasets in the United States, our study will support communities across Minnesota and the United States in their ongoing efforts to improve road safety for pedestrians and be helpful for automated vehicle design.
@@ 1 Introduction
2 Review of relevant literature
The section is divided into four main sections: the first identifies driver characteristics, the second discusses pedestrian characteristics and behavior, the third focuses on traffic and vehicle characteristics, and the fourth addresses intersection characteristics related factors influencing driver yielding behavior towards pedestrians, as identified in previous literature. After these four sections, we include another section where we mention the gap that our study aims to address.
2.1 Driver characteristics
A number of studies have focused on driver characteristics related factors influencing driver behavior at intersections, encompassing both signalized and unsignalized types, as well as mid-block crossings [11, 12, 13, 8]. This research has explored the role of demographic factors such as age and educational background on driver-yielding behavior. [14] observed that older drivers are more prone to yield to pedestrians than their younger counterparts. The educational level of drivers also plays a significant role, with those holding a bachelor’s degree demonstrating a higher tendency to yield compared to drivers with lower educational levels [14]. This trend suggests that awareness and comprehension of traffic regulations, including right-of-way rules and the concept of unmarked crosswalks, are instrumental in guiding driver behavior [4].
Further emphasizing the influence of awareness campaigns, [15] noted the effectiveness of educational campaigns in positively shaping driver-yielding practices. A notable example of this is a study by [16], which reported a significant decrease in non-yielding behavior, from 61.42% to 55.19%, following a targeted campaign on a university campus.
One significant finding is that driver-yielding behavior is further influenced by the presence of a following vehicle [17, 18] and the actions of drivers in the opposite direction [6, 17], highlighting the complexity of factors at play.
2.2 Pedestrian characteristics and behavior
@@
In addition to driver characteristics, pedestrian characteristics and behavior at intersections can influence the drivers’ yielding behavior [19, 20, 21, 8]. For example, there is a methodology for analyzing unsignalized intersections based on gap acceptance behavior [22], and gap acceptance theory is widely used to evaluate the safe operation of the road at uncontrolled intersections [19]. Pedestrians have to wait for safe gaps to cross at intersections. This is a high risk for pedestrians because the wrong judgment of the accepted gap can result in traffic crashes. From the perspective of a pedestrian, a pedestrian will either accept or reject the gap in the situation of vehicle-pedestrian interaction, and this will force the driver to react differently if the pedestrian accepts a shorter gap.
Studies have shown that drivers are more inclined to yield to older pedestrians than younger ones [23]. A study by [8] shows that drivers demonstrated a higher tendency to yield to pedestrians who were female, crossing in a group, carrying baggage, not using a mobile phone, making some hand gesture to the driver, and employing the rolling gap strategy while crossing. In addition to that pedestrian assertiveness, such as briskly walking towards the crosswalk, is also a key factor leading to higher-yielding rates [17, 23, 24].
Interestingly, visible cues significantly influence driver behavior. Research indicates that pedestrians using a white cane or donning bright clothing are more likely to elicit yielding from drivers, in contrast to those holding an umbrella or wearing less conspicuous attire [25, 26]. Additionally, the positioning of pedestrians at the crosswalk plays a critical role; those positioned less than one foot from the curb are observed to have a higher likelihood of being yielded to [27].
Drivers are less likely to yield to pedestrians outside marked crosswalks, and they tend to decelerate more for pedestrians within a crosswalk than for those outside it. Collectively, these insights contribute to a nuanced understanding of pedestrian dynamics at intersections, highlighting areas for potential intervention and planning.
2.3 Vehicle and traffic characteristics
The influence of vehicle characteristics on driver-yielding behavior is a subject of considerable research interest. A notable finding is the tendency of drivers in high-status, expensive vehicles in the United States to yield less frequently to pedestrians, indicating a relationship between vehicle status and driver behavior [28].
In addition to vehicle status, speed is a critical factor that significantly impacts driver yielding. Research has consistently shown that lower vehicle speeds are associated with higher yielding rates [29, 17, 26, 30, 31]. [27] observed this trend in a study focusing on a two-lane roundabout. They reported that a decrease in vehicle speed to 15 miles per hour resulted in a yielding rate increase of up to 75%. In contrast, an increment of merely five miles per hour led to a notable reduction in yielding rates, dropping them to about 50%. Furthermore, specific vehicle characteristics, such as model and type, also play a significant role in influencing driver behavior at intersections [18].
Beyond individual vehicle attributes, the overall traffic dynamics, such as vehicular volume and the presence of vehicle platoons, crucially affect driver-yielding behavior. Studies have specifically highlighted the impact of vehicle platoons, showing that they tend to decrease the likelihood of drivers yielding to pedestrians [17, 32].
2.4 Intersection characteristics
The design and layout of intersections play a pivotal role in driver behavior towards pedestrians. This aspect of road safety, extending beyond individual driver or pedestrian behavior, has been the focus of various studies. The idea is that certain intersection characteristics may enhance or compromise pedestrian safety, influencing how welcoming these spaces are for pedestrians.
Research has consistently shown that the physical layout of roads and their surrounding environments significantly impacts driver-yielding behavior. One key finding is that drivers tend to yield less frequently on roads where pedestrians face longer crossing distances [33, 34]. [24] reported lower yielding rates on major roadways, such as arterial and collector, particularly at unmarked crosswalks, compared to local roads with marked crosswalks.
The angle of intersection is another crucial factor. Studies indicate that left-turning drivers’ yielding behavior is influenced by the intersection angle. Drivers at obtuse-angled intersections tend to exhibit more non-yielding behavior, whereas those at acute-angled intersections are more likely to yield [35, 32]. Moreover, the geometric characteristics of intersections also affect vehicle speed and the potential for collisions.
Effective engineering treatments, including signalization, crossing marks, red signal or beacon treatments, and pedestrian-activated signals, have been shown to significantly influence yielding behavior [36, 37]. Additionally, road signs prompting drivers to yield and pedestrian lights, though their activation rates may vary with traffic conditions, have been effective in increasing safe yielding distances [38].
Further emphasizing the importance of design, features such as marked pedestrian crossings, high-visibility crosswalks, overhead warning signs, rectangular rapid flashing beacons, advance yield signs, and distinct road markings have been identified as effective in enhancing pedestrian safety. These elements encourage drivers to yield more readily and are critical in guiding the construction of new intersections or the modification of existing ones to improve pedestrian safety [15, 34].
2.5 Study gap
Based on the literature reviewed, several features related to drivers, pedestrians, traffic, vehicles, and intersection characteristics lead to higher driver-yielding rates. However, previous research does not consider the effects of the surrounding built environment on driver-yielding behavior towards pedestrians. Numerous built environment factors, such as the presence of parking, bike lanes, schools, varying land use types, and bus stops, have been recognized for their influence on pedestrian crashes and their severity level [39, 40, 41, 42]. Despite these built environmental elements having the potential to significantly affect how drivers respond to pedestrians, detailed investigations into the relationship between these site-specific built environment features and driver yielding are sparse. This research gap is what our study aims to address, utilizing large-scale open-source data to explore the potential influence of built environment factors on driver-yielding behavior in a comprehensive manner.
@@ 3 Data Collection and Description
One of the contributions of this work is the collection of a large-scale, open-source naturalistic dataset for driver yielding. We collected over 3,300 driver-pedestrian interactions at 18 different unsignalized intersections. The dataset, which focuses on driver-pedestrian interactions, is unique in both its scale and the depth of information it provides. In this section, we first describe site characteristics and the data collection technique, followed by a summary of the selected sites, and finally present summary statistics of the collected data. It is important to note that the sites are anonymized in our reporting to maintain privacy standards.
@@ 3.1 Site Characteristics
The selection of sites was guided by a set of criteria derived from the literature reviewed in Section 2. These criteria include factors like traffic volume, pedestrian density, and urban setting, as identified by Minnesota Department of Transportation (MnDOT) staff. Our chosen sites, spread across diverse geographical locations, provide a representative cross-section of various urban traffic scenarios and were selected with the guidance of MnDOT staff.
The table LABEL:tab:sites below summarizes the characteristics of each site, including the number of lanes, posted speed limits, type of road markings, AADT, and intersection geometry. This diversity in site characteristics allows for a robust analysis of driver-pedestrian interactions under different traffic conditions.
Site Num. of Lanes (Main Street) Num. of Lanes (Minor Street) Posted Speed (MPH) Markings AADT Shape Site 1 4 2 35 Unmarked 14600 T-shape Site 2 2 2 35 Unmarked 14800 T-shape Site 3 2 1 30 Standard1 14800 T-shape Site 4 2 1 30 Unmarked 10700 four-way Site 5 2 1 30 Unmarked 10500 four-way Site 6 3 1 35 Continental 6200 four-way Site 7 4 1 35 Continental 3300 T-shape Site 8 4 1 30 Unmarked 10200 four-way Site 9 3 1 35 Unmarked 5400 four-way Site 10 4 1 35 Unmarked 18800 four-way Site 11 3 1 30 Unmarked 18100 four-way Site 12 2 1 30 Unmarked 10300 four-way Site 13 2 1 35 Continental 7400 four-way Site 14 2 1 30 Standard 7100 four-way Site 15 3 1 30 Unmarked 16400 four-way Site 16 2 1 30 Standard 12100 four-way Site 17 4 1 30 Unmarked 12800 T-shape Site 18 4 1 30 Unmarked 12500 four-way 1 Standard markings refer to markings with two solid white lines.
@@ 3.2 Data Collection and Cleaning
To investigate driver-yielding behavior, we utilized the traffic information monitor (TIM) platform developed by the Minnesota Traffic Observatory (MTO) at the University of Minnesota. The TIM is a customized sensor pack that consists of a pole-mounted video camera that can extend up to 30 feet, 12 V batteries that can sustain up to four weeks of operation, a Raspberry Pi computer, and a watertight metal enclosure to protect the batteries and computer. In this study, we used the TIM and the camera to capture video data of pedestrians and drivers approaching intersections in a naturalistic setting, as depicted in Fig. 1.
@@
In the summer of 2021, we strategically deployed TIMs across 18 diverse sites in Minnesota. These sites were chosen to encapsulate a broad spectrum of pedestrian and driver interactions, environmental contexts, and traffic patterns. The TIMs recorded extensive video data during daylight hours for one to two weeks at each site, focusing on capturing naturalistic driver and pedestrian behavior at intersections.
@@
We collected more than 50 features for each interaction, and each pair of interactions needs to be collected individually. To extract the driver-pedestrian interactions accurately and analyze the data efficiently, a trained data collector manually extracted individual interaction events of all the variables. The speed extraction methodology computes the vehicular speed of the concerned driver-pedestrian interaction by measuring the time each vehicle takes to pass two known positions at the intersection. Building upon preliminary work cited in [10], which involved data from three sites, this study expands the dataset to include 18 locations.
@@
In this study, a “codable driver-pedestrian interaction” is characterized as an event where either the driver yields to the pedestrian or the pedestrian forces to yield for giving the right of way to the vehicle. A “close call” is defined as a near-miss incident in which an imminent collision between a vehicle and a pedestrian is narrowly averted, often necessitating quick evasive actions by either or both parties involved. It’s worth noting that this measure is somewhat subjective, especially given that data collection was performed by a single individual. The term “number of bus stops” refers to the count of bus stops located within one block of the main road intersecting the observed location. For further details on the variables collected and their implications, we direct readers to our publicly available repository of all the collected data and Tables LABEL:tab:sites and 2 in the following sections, and Table 5 in the Appendix. We acknowledge the potential subjectivity in data collection. To mitigate this, we provided a clear definition of the variables during the data collection procedure. Additionally, having one consistent data collector, who worked on this project for over a year, helped ensure uniformity in the data collection process.
3.3 Data Description
Table 2 presents the summary statistics for pedestrian behavior at the 18 intersections studied. Factors include the number of total interactions, the number of close calls, and the driver’s yielding rate. Intersection 16 had the most interactions with 840 events, and intersection 11 had the fewest number of interactions with 20. While intersection 16 also had the highest yielding rate of 70.36%, intersection 11 had the lowest yielding rate of 0. This suggests that locations with higher pedestrian volumes tend to have higher driver-yielding rates. Overall, we found that yielding rates are generally low, as shown in Table 2. For example, at Site 5, the yielding rate was particularly low, and at intersections 11 and 15, the yielding rate was 0, indicating that no drivers yielded to pedestrians during our study period. This is likely due to the environmental factors at these sites, where drivers may not anticipate yielding to pedestrians. Another noteworthy observation is that the maximum waiting time for pedestrians at certain intersections approaches 1 minute. In some instances, pedestrians were forced to halt midway through crossing, potentially elevating the risk of crashes. The diminished yielding rate at unsignalized, unmarked intersections with low pedestrian volume may be attributed to drivers’ lack of anticipation for pedestrian presence.